### v1

# Synthesizing High-Quality Visual Question Answering from Medical Documents with Generator-Verifier LMMs

Abstract
Large Multimodal Models (LMMs) are increasingly capable of answering medical questions that require joint reasoning over images and text, yet training general medical VQA systems is impeded by the lack of large, openly usable, high-quality corpora. We present MedVLSynther, a rubric-guided generator-verifier framework that synthesizes high-quality multiple-choice VQA items directly from open biomedical literature by conditioning on figures, captions, and in-text references. The generator produces self-contained stems and parallel, mutually exclusive options under a machine-checkable JSON schema; a multi-stage verifier enforces essential gates (self-containment, single correct answer, clinical validity, image-text consistency), awards fine-grained positive points, and penalizes common failure modes before acceptance. Applying this pipeline to PubMed Central yields MedSynVQA: 13,087 audited questions over 14,803 images spanning 13 imaging modalities and 28 anatomical regions. Training open-weight LMMs with reinforcement learning using verifiable rewards improves accuracy across six medical VQA benchmarks, achieving averages of 55.85 (3B) and 58.15 (7B), with up to 77.57 on VQA-RAD and 67.76 on PathVQA, outperforming strong medical LMMs. Ablations verify that both generation and verification are necessary and that more verified data consistently helps, and a targeted contamination analysis detects no leakage from evaluation suites. By operating entirely on open literature and open-weight models, MedVLSynther offers an auditable, reproducible, and privacy-preserving path to scalable medical VQA training data.
1 Introduction
Large Multimodal Models (LMMs) are rapidly becoming capable assistants for biomedical discovery and clinical education, where questions must be answered by jointly interpreting medical images (e.g., X‑ray, CT, microscopy) and the surrounding textual context (e.g., figure captions, narrative descriptions, etc.). Despite fast progress, training general medical VQA systems remains difficult because the community lacks large, openly usable, and high‑quality training corpora.
On the evaluation side, recent benchmark (Hu et al., 2024; Ye et al., 2024) provide broad and challenging test suites, but they are designed for assessment rather than training and therefore offer no training splits. On the training side, existing datasets fall into three categories, each with a limitation. 1) Manually curated sets (Lau et al., 2018; Liu et al., 2021; He et al., 2020) are carefully annotated but are either small or bound to narrow modalities and topics, limiting coverage and transfer. 2) Automatically generated sets (Zhang et al., 2023b; Chen et al., 2024c) scale more easily but are typically produced by text‑only LLMs that ignore visual evidence and figure–text relations, yielding noisy stems, ambiguous options, and medically dubious answers that can impede model learning. 3) Closed, large‑scale resources (Li et al., 2024) exist but are not publicly shareable due to patient privacy, licensing, and institutional agreements, which slows open research and reproducibility. Collectively, these constraints lead to a practical bottleneck: we can evaluate medical VQA systems comprehensively, but we cannot train them broadly and transparently.
This paper asks a simple question: can we synthesize high‑quality, auditable medical VQA data directly from open biomedical literature? Our answer is MedVLSynther, a generator–verifier framework that leverages state‑of‑the‑art open‑weight LMMs (Zeng et al., 2025; Wang et al., 2025; Bai et al., 2025) to produce and automatically vet VQA triplets from figures and surrounding text in PubMed articles (Lozano et al., 2025; Roberts, 2001). The key design choice is to make both generation and verification explicitly rubric‑driven and context‑aware.
Rubric‑guided context‑aware generation (Figure 1 (a)). Given a figure, its caption, and the figure’s in‑text reference paragraph when available, the generator LMM is instructed to propose a VQA item, including question stem, multiple‑choice options, and the correct answer, under a comprehensive rubric. The rubric enforces that stems are self‑contained and anchored in the provided visual–textual context, that options are parallel and mutually exclusive, and that the answer can be justified from the figure and caption, not from world knowledge alone. The rubric also specifies a set of accepted question archetypes (e.g., recognition, localization, comparative, reasoning) and a JSON schema/format that simplifies downstream filtering and training.
Multi‑stage rubric‑based verification (Figure 1 (b)). To ensure quality, we feed the same context and the generated VQA to a verifier LMM and score it in three stages: 1) Essential criteria form strict pass/fail gates. Any failure discards the item. 2) Fine‑grained criteria award positive points with justifications, and allow the verifier to surface additional criteria opportunistically. 3) Penalty criteria investigate common failure modes and subtract points when detected. We sum the fine‑grained and penalty scores and apply a threshold to filter surviving items. This verifier is model‑agnostic and can be instantiated with any open‑weight LMM; in practice we find that a verifier different from the generator improves robustness.
The generator–verifier loop yields a data pipeline whose rules are transparent and auditable end‑to‑end. Because we build on open literature rather than protected clinical data, the entire pipeline, including prompts, rubric, and metadata, can be inspected and reproduced. At the same time, recent open‑weight LMMs rival proprietary systems on many multimodal tasks (Zeng et al., 2025), allowing us to benefit from strong perception and reasoning while staying fully open.
The resulting medical VQA dataset, MedSynVQA, covering diverse modalities, subspecialties, and question archetypes. Models trained on this data with Reinforcement Learning with Verifiable Rewards (RLVR) (Guo et al., 2025; Shao et al., 2024) outperform counterparts trained on PMC‑VQA (Zhang et al., 2023b), as well as the strong baseline trained on text‑only medical corpora (Huang et al., 2025b). As summarized in Figure 1 (c), our training improves accuracy on MMMU‑Med (Yue et al., 2024), PathVQA (He et al., 2020), and VQA‑RAD (Lau et al., 2018) over strong baselines (Alshibli et al., 2025; Li et al., 2023). Meanwhile, ablations reveal that (i) both generation and verification are necessary: their synergy yields the best accuracy, and (ii) scale matters: more verified data consistently helps. We analyze topic coverage, modality distribution, and question types, and most importantly, conduct a contamination analysis tailored for synthetic medical VQA; we find no detectable leakage from the evaluation sets.
Our contributions are summarized as follows:
MedVLSynther, a rubric‑guided, context‑aware generator–verifier pipeline that synthesizes reliable medical VQA from open biomedical articles.
A comprehensive rubric for medical VQA quality, spanning essential gates, fine‑grained positive criteria, and penalty criteria, together with a machine‑checkable schema that supports automatic filtering and auditing.
A synthetic medical VQA training set (MedSynVQA) that substantially improves medical LMMs on multiple medical VQA benchmarks and complements existing resources without relying on private patient data.
Transparency and reproducibility: our pipeline operates entirely on open data and open models, enabling the community to inspect prompts, scoring rules, and filtering decisions end‑to‑end.
While synthetic data cannot replace carefully curated clinical datasets, our results indicate that high‑quality, auditable synthesis is both feasible and useful for medical VQA. We hope MedVLSynther provides a practical path to scalable training data that respects privacy, encourages openness, and accelerates progress in multimodal medical intelligence.
| Dataset | # Questions | # Images | # Rate | # Modality | # Anatomy | Annotation | Data Availability | General QA | Training Set |
| MedXpertQA-MM | 2,000 | 2,852 | 1.43 | 10 | 11 | Expert | Open access | Yes | No |
| GMAI-MMBench | 25,831 | 25,831 | 1.00 | 38 | N/A | Automatic | Open access | Yes | No |
| OmniMedVQA | 127,995 | 118,010 | 0.92 | 12 | 26 | Automatic | Open access | Yes | No |
| SLAKE | 14,028 | 642 | 0.05 | 3 | 5 | Expert | Open access | No | Yes |
| VQA-RAD | 3,515 | 315 | 0.09 | 3 | 3 | Expert | Open access | No | Yes |
| PathVQA | 32,799 | 4,998 | 0.15 | 2 | N/A | Automatic | Open access | No | Yes |
| PMC-VQA | 226,946 | 149,075 | 0.66 | N/A | N/A | Automatic | Open access | Yes | Yes |
| GMAI-VL-5.5M | 5,500,000 | N/A | N/A | 13 | N/A | Automatic | Not Open | Yes | Yes |
| MedSynVQA | 13,087 | 14,803 | 1.13 | 13 | 28 | Automatic | Open access | Yes | Yes |
2 Related Works
Multimodal medical VQA.
Early, expert‑curated datasets (Lau et al., 2018; Liu et al., 2021; He et al., 2020) established Med‑VQA but remain small or modality‑restricted, limiting transfer. Later, broad benchmarks (Hu et al., 2024; Ye et al., 2024; Zuo et al., 2025) consolidated evaluation across many modalities and anatomies yet offer little or no training data, creating a supervision bottleneck. In contrast, large literature‑derived corpora (Subramanian et al., 2020; Rückert et al., 2024) and especially Lozano et al. (2025)’s 24M image–caption pairs provide open, scalable raw material. Our work converts this open substrate into exam‑quality VQA by coupling context‑aware generation with rigorous verification, bridging the gap between expansive evaluation suites and accessible training data.
Synthetic data generation for multimodal medical VQA.
Prior synthetic pipelines scale supervision but suffer quality issues: Li et al. (2023)’s self‑instruct approach and Zhang et al. (2023b)’s 227k auto‑generated pairs (largely from text‑only LLMs) can omit modality cues, produce ambiguous stems, and yield visually ungrounded answers; broader compilations like Li et al. (2024) are closed, while modality‑specific (Hu et al., ) sets remain narrow. These limitations motivate a quality‑first strategy: we condition on figures, captions, and in‑text references and enforce a rubric‑guided generator plus a multi‑stage verifier to filter low‑quality items, yielding reliable, open data suitable for training medical LMMs without relying on private images.
Multimodal models, medical adaptation, and reasoning.
General LVLMs (Hurst et al., 2024; Comanici et al., 2025; Wang et al., 2025; Bai et al., 2025; Liu et al., 2024; An et al., 2025; ) acquire instruction following via visual SFT, while medical variants (Tu et al., 2024; Luo et al., 2023; Alshibli et al., 2025; Liu et al., 2023; Wu et al., 2025; Chen et al., 2024b; Zhou et al., 2024; Wu et al., 2023) add in‑domain pretraining and SFT/RL for clinical competence. Recent deliberate‑reasoning models (Jaech et al., 2024) show that reinforcement learning with verifiable rewards (e.g., GRPO (Guo et al., 2025)) can surpass SFT‑only methods on multi‑step problems, and early medical efforts point the same way but lack open, high‑quality multimodal supervision. Our rubric‑verified VQA corpus supplies that missing signal and pairs naturally work for RLVR, contributing auditable and trustworthy visual reasoning in open‑weight medical LMMs (Chen et al., 2024c; Su et al., 2025).
3 MedVLSynther and MedSynVQA
Our goal is to synthesize high‑quality, clinically valid multiple‑choice VQA (MC‑VQA) examples directly from biomedical papers (Lozano et al., 2025). We cast the task as a Generator–Verifier pipeline driven by Large Multimodal Models (LMMs): a rubric‑guided generator produces MC‑VQA items from figures and text, and a multi‑stage rubric‑guided verifier performs automatic quality control before data are admitted to the final corpus (Figure 3).
3.1 Data source, extraction, and pre‑filtering
Source.
We build on Biomedica (Lozano et al., 2025), a large‑scale extraction of figures and figure‑level metadata from the PubMed Central Open‑Access (PMC-OA) collection (Roberts, 2001). For each paper we ingest: 1) The figure image(s) (a single caption may reference up to 6 images), 2) The figure caption. 3) The corresponding figure references in the main text (when present).
Samples missing either images or a caption are discarded.
Pre‑filtering.
We retain items annotated by Lozano et al. (2025) with the primary labels: Clinical imaging and Microscopy, and 25 secondary subtypes (e.g., x‑ray radiography, optical coherence tomography, skull, brain, etc.). After pre‑filtering we obtain 23,788 figure‑caption(-reference) triplets.
We denote each pre‑filtered sample by
where is one or more images, the caption, and the in‑text references.
Choice of generator and verifier LMMs.
We use state‑of‑the‑art open‑weight LMM capable of long‑context vision‑language reasoning: GLM‑4.5V‑108B (Zeng et al., 2025), InterVL‑3.5‑38B (Wang et al., 2025), and Qwen2.5‑VL‑72B (Bai et al., 2025). Unless otherwise noted, GLM‑4.5V‑108B serves as the default generator due to its strong instruction‑following and image‑grounding performance. The rubric and strict JSON schema make the output predictable and machine‑verifiable.
3.2 Rubric‑based, context‑aware VQA generation
Given , the generator LMM produces a 5‑option MC‑VQA instance in strict JSON format: . Generation is context‑aware the model receives the image(s) together with and . To ensure exam‑quality items, the prompt instills the role of an expert medical‑education item writer and enforces a self‑check rubric.
Essential (must pass before output): 1) Stem self‑contained (no “caption/context” mentions); 2) Image–content alignment (requires inspecting specific visual features); 3) Implicit use of caption facts without answer leakage; 4) Exactly one best answer; 5) Medical correctness (modality, anatomy, terminology).
Important (strongly recommended): cognitive level is over application; strong, parallel distractors; clear focus on a single concept.
Optional: localization or quantitative details when clearly supported.
A small set of question archetypes (i.e., finding identification, diagnosis, next step, localization, modality recognition) reduces prompt entropy and encourages clinically meaningful questions.
3.3 Multi‑stage, rubric‑based, context‑aware verification
While the generator is reliable, automatic verification is essential for scale and precision. Given and a candidate MC‑VQA , the verifier LMM is prompted to operate in two roles, Referee and Critic, and to return only a structured rubric with binary scores. Verification is also context‑aware: sees the same images, caption, and references as plus the proposed MC‑VQA.
Stage‑1: Essential screening (hard gate).
The Referee evaluates seven non‑negotiable items; a sample must pass all to proceed: 1) Stem Self‑contained; 2) Vocabulary Constraint (no unsupported clinical facts); 3) Diagnosis Leak (no verbatim restatement from sources); 4) Single Correct Option; 5) Option Type Consistency (same semantic type); 6) Clinical Validity (terminology/modality/anatomy); 7) Image–Text Consistency.
Items are scored with a fair, rule‑based mindset. During this stage, we remove instances that the verifier cannot grade (e.g., malformed JSON), leaving 23,635 candidates. After applying the essential filter, 22,903 remain.
Stage‑2: Fine‑grained positive criteria (bonus points).
The Critic now assumes the item is not excellent and awards points only on irrefutable evidence We query 4–8 bonus criteria (binary, with weights Important or , Optional or ), including: 1) Plausible Distractors (every distractor is a strong near‑miss); 2) Parallel Options (length/structure uniformity); 3) Stem Concision (less than two sentences and concise); 4) Clarity and Focus (single, unambiguous question); 5) Answer‑field Validity (answer exists, matches an option); 6) JSON Schema Compliance (exact keys, no extras).
The Critic denies a criterion if it can imagine a slightly better wording or distractor, pushing precision over recall.
Stage‑3: Penalty criteria (error hunting).
Finally, the Critic actively searches for pitfalls (negative weights): 1) Forbidden Terms (; stem contains “caption/context”); 2) Synonym Drift (; introduces unsupported specific facts); 3) Multiple Keys (), and Medical Inaccuracy ().
Each pitfall is triggered only with a concrete reason.
3.4 Aggregation and acceptance rule
Let be the set of positive (Important Optional) criteria with weights and binary scores . Let be the pitfalls with and scores . We compute a normalized quality score:
Candidates passing Stage‑1 are accepted if with . This high threshold emphasizes precision while keeping a useful yield; it results in 13,087 MC‑VQA items, which we call MedSynVQA.
3.5 Training medical LMMs with MedSynVQA
We use our synthesized corpus to train medical LMMs with two LMM finetuning approaches.
Supervised Fine‑Tuning (SFT). Following MedVLThinker, we elicit thinking traces with GLM‑4.5V‑108B and perform SFT on (thinking trace, answer) pairs. The supervision emphasizes clinically grounded reasoning paths while preserving the strict answer format.
RL with Verbal Rewards (RLVR). We then apply GRPO on answers only (no trace optimization), again mirroring hyper‑parameters from Huang et al. (2025b). The reward promotes exact‑match accuracy and adherence to the schema without over‑fitting to any single imaging modality.
| Model | MMMU | MedX-M | PathVQA | PMC | SLAKE | VQA-RAD | Avg. |
| Qwen2.5-VL-3B-Instruct | 44.12 | 20.69 | 61.96 | 44.77 | 61.30 | 62.01 | 49.14 |
| PMC-Style Text-only Generation | 48.82 | 20.40 | 63.38 | 51.08 | 73.08 | 72.06 | 54.80 |
| Rubric Context-Aware Generation | 52.35 | 20.60 | 62.49 | 51.83 | 70.43 | 70.59 | 54.72 |
| + Rubric Context-Aware Verification | 52.35 | 21.40 | 62.82 | 50.23 | 74.76 | 73.53 | 55.85 |
| Qwen2.5-VL-7B-Instruct | 52.94 | 18.89 | 65.39 | 49.30 | 65.71 | 68.75 | 53.50 |
| PMC-Style Text-only Generation | 51.76 | 21.70 | 64.31 | 53.43 | 68.03 | 71.69 | 55.15 |
| Rubric Context-Aware Generation | 58.24 | 23.50 | 65.41 | 53.83 | 68.03 | 75.00 | 57.33 |
| + Rubric Context-Aware Verification | 57.06 | 23.15 | 66.36 | 53.78 | 67.79 | 77.21 | 57.56 |
| Model | Scale | MMMU | MedX-M | PathVQA | PMC | SLAKE | VQA-Rad | Avg. |
| Qwen2.5-VL 3B-Instruct | N/A | 44.12 | 20.69 | 61.96 | 44.77 | 61.30 | 62.01 | 49.14 |
| 1000 | 50.59 | 20.20 | 63.18 | 48.37 | 65.87 | 67.65 | 52.64 | |
| 2000 | 47.06 | 19.95 | 64.07 | 47.27 | 74.04 | 76.84 | 54.87 | |
| 5000 | 52.35 | 21.40 | 62.82 | 50.23 | 74.76 | 73.53 | 55.85 | |
| 10000 | 48.82 | 20.55 | 63.44 | 49.87 | 72.84 | 74.63 | 55.03 | |
| Full | 51.76 | 22.30 | 63.03 | 48.92 | 72.60 | 72.43 | 55.17 | |
| Qwen2.5-VL 7B-Instruct | N/A | 52.94 | 18.89 | 65.39 | 49.30 | 65.71 | 68.75 | 53.50 |
| 1000 | 57.65 | 21.60 | 65.53 | 50.93 | 68.27 | 65.81 | 54.96 | |
| 2000 | 60.00 | 22.35 | 67.76 | 51.18 | 67.31 | 73.16 | 56.96 | |
| 5000 | 57.06 | 23.15 | 66.36 | 53.78 | 67.79 | 77.21 | 57.56 | |
| 10000 | 57.06 | 22.45 | 66.86 | 52.73 | 71.88 | 73.90 | 57.48 | |
| Full | 55.88 | 22.10 | 65.56 | 55.43 | 72.36 | 77.57 | 58.15 |
| Model | Generator | Verifier | MMMU | MedX-M | PathVQA | PMC | SLAKE | VQA-Rad | Avg. |
| Qwen2.5-VL 3B-Instruct | N/A | N/A | 44.12 | 20.69 | 61.96 | 44.77 | 61.30 | 62.01 | 49.14 |
| GLM-4.5V 108B | Qwen2.5-VL 72B | 52.35 | 21.40 | 62.82 | 50.23 | 74.76 | 73.53 | 55.85 | |
| GLM-4.5V 108B | GLM-4.5V 108B | 51.18 | 20.30 | 63.56 | 50.63 | 71.63 | 70.22 | 54.59 | |
| Qwen2.5-VL 72B | GLM-4.5V 108B | 47.65 | 21.50 | 62.37 | 48.87 | 73.32 | 69.85 | 53.93 | |
| InternVL3.5 38B | GLM-4.5V 108B | 49.41 | 21.90 | 61.81 | 51.98 | 74.76 | 71.32 | 55.20 | |
| Qwen2.5-VL 7B-Instruct | N/A | N/A | 52.94 | 18.89 | 65.39 | 49.30 | 65.71 | 68.75 | 53.50 |
| GLM-4.5V 108B | Qwen2.5-VL 72B | 57.06 | 23.15 | 66.36 | 53.78 | 67.79 | 77.21 | 57.56 | |
| GLM-4.5V 108B | GLM-4.5V 108B | 58.82 | 23.65 | 67.22 | 54.48 | 71.15 | 73.16 | 58.08 | |
| Qwen2.5-VL 72B | GLM-4.5V 108B | 56.47 | 22.55 | 67.25 | 52.38 | 67.07 | 72.79 | 56.42 | |
| InternVL3.5 38B | GLM-4.5V 108B | 57.65 | 23.30 | 66.12 | 53.58 | 70.67 | 75.37 | 57.78 |
| Model | MMMU | MedX-M | PathVQA | PMC | SLAKE | VQA-RAD | Avg. |
| Qwen2.5-VL-3B-Instruct | 44.12 | 20.69 | 61.96 | 44.77 | 61.30 | 62.01 | 49.14 |
| SFT (PMC) | 47.84 | 21.46 | 52.76 | 54.55 | 65.79 | 58.58 | 50.16 |
| SFT (m23k) | 32.55 | 16.00 | 42.74 | 28.53 | 43.91 | 33.09 | 32.80 |
| SFT (MedSynVQA) | 48.82 | 20.90 | 63.12 | 47.57 | 54.33 | 59.93 | 49.11 |
| RL (PMC) | 48.43 | 21.51 | 51.61 | 54.22 | 75.56 | 62.38 | 52.28 |
| RL (m23k) | 52.16 | 22.90 | 62.28 | 47.32 | 63.38 | 71.08 | 53.19 |
| RL (MedSynVQA) | 52.35 | 21.40 | 62.82 | 50.23 | 74.76 | 73.53 | 55.85 |
| Qwen2.5-VL-7B-Instruct | 52.94 | 18.89 | 65.39 | 49.30 | 65.71 | 68.75 | 53.50 |
| SFT (PMC) | 49.80 | 21.39 | 53.02 | 54.67 | 67.71 | 57.72 | 50.72 |
| SFT (m23k) | 46.86 | 16.40 | 56.35 | 34.58 | 54.97 | 53.80 | 43.83 |
| SFT (MedSynVQA) | 49.41 | 20.90 | 64.81 | 50.08 | 59.62 | 66.54 | 51.89 |
| RL (PMC) | 55.29 | 24.11 | 57.09 | 55.38 | 66.59 | 63.48 | 53.66 |
| RL (m23k) | 56.86 | 24.43 | 66.83 | 50.67 | 65.79 | 64.71 | 54.88 |
| RL (MedSynVQA) | 57.06 | 23.15 | 66.36 | 53.78 | 67.79 | 77.21 | 57.56 |
| Model | MMMU | MedX-M | PathVQA | PMC | SLAKE | VQA-Rad | Avg. |
| General LLM | |||||||
| Gemme 3 4B | 46.67 | 21.89 | 59.24 | 44.42 | 66.59 | 56.86 | 49.28 |
| Qwen2.5-VL-3B-Instruct | 44.12 | 20.69 | 61.96 | 44.77 | 61.30 | 62.01 | 49.14 |
| Qwen2.5-VL-7B-Instruct | 52.94 | 18.89 | 65.39 | 49.30 | 65.71 | 68.75 | 53.50 |
| Medical LLM | |||||||
| MedGemma 4B | 32.55 | 8.17 | 59.64 | 42.73 | 83.49 | 78.55 | 50.86 |
| MedGemma 27B | 35.88 | 12.13 | 62.09 | 36.75 | 77.40 | 72.67 | 49.49 |
| Llava Med V1.5 7B | 31.37 | 22.56 | 56.52 | 34.28 | 62.82 | 56.74 | 44.05 |
| HuatuoGPT-Vision-7B | 50.59 | 22.00 | 63.53 | 53.39 | 75.00 | 63.60 | 54.69 |
| MedVLThinker-3B | 52.16 | 22.90 | 62.28 | 47.32 | 63.38 | 71.08 | 53.19 |
| MedVLThinker-7B | 56.86 | 24.43 | 66.83 | 50.67 | 65.79 | 64.71 | 54.88 |
| MedVLSynther-3B | 52.35 | 21.40 | 62.82 | 50.23 | 74.76 | 73.53 | 55.85 |
| MedVLSynther-7B | 55.88 | 22.10 | 65.56 | 55.43 | 72.36 | 77.57 | 58.15 |
4 Experiments
4.1 Setup
Models.
Unless otherwise stated, we finetune two open‑weight LMMs Qwen2.5‑VL 3B and 7B Instruct (Bai et al., 2025), using the same training schedule, image resolution, tokenization, and optimization hyper‑parameters as Huang et al. (2025b). We use our rubric‑guided generator–verifier pipeline to synthesize training items from PubMed figures and captions (Figure 3), and we train students either with SFT or RLVR. Unless otherwise noted, experiments use 5K samples.
Benchmarks and metric.
We follow Huang et al. (2025b) evaluation suite and scripts, reporting multiple‑choice accuracy on six medical VQA benchmarks: MMMU medical split (MMMU-Med) (Yue et al., 2024), (MedX‑M) (Zuo et al., 2025), PathVQA (He et al., 2020), PMC‑VQA (Zhang et al., 2023b), SLAKE (Liu et al., 2021), and VQA‑RAD (Lau et al., 2018).
Baselines.
We compare against strong general‑purpose and medical LMMs used in Huang et al. (2025b), including Gemma3 4B (Team et al., 2025), Qwen2.5‑VL‑3B/7B‑Instruct, MedGemma 4B (Sellergren et al., 2025), LLaVA‑Med (Li et al., 2023), HuatouGPT‑Vision‑7B (Chen et al., 2024c), and MedVLThinker (Huang et al., 2025b), strong baselines trained solely on text-only data.
4.2 Results
Ablation on the Generator–Verifier pipeline.
Table 2 studies each stage of our pipeline. We begin from zero‑shot Qwen2.5‑VL students and add 1) PMC‑style text‑only question generation, 2) rubric‑guided context‑aware generation, and 3) rubric‑aware verification. For 3B student, the base model averages 49.14. Text‑only generation lifts the average to 54.80. Switching to rubric‑guided, context‑aware generation performs similarly on average (54.72). Adding verification yields the best average, 55.85, with large gains on clinically grounded datasets. For 7B student, the base model averages 53.50. Text‑only generation yields 55.15, rubric‑guided generation 57.33, and with verification we obtain 57.56 and again improving across benchmarks. Overall, rubric guidance already outperforms a PMC‑style text‑only recipe, and multi‑stage verification supplies the remaining headroom, producing the best average in both scales (Table 2). The trend aligns with the high‑level improvements visualized in Figure 1.
How much synthesized data do we need?
We vary the number of MedSynVQA training items from 1K to 13K (Table 3). For 3B student. Accuracy increases from 52.64 (1K) to 55.85 (5K), then plateaus at 55.03 (10K) and 55.17 (Full). For 7B student. The curve is similar: 54.96 (1K), 56.96 (2K), 57.56 (5K) with a slight dip to 57.48 at 10K, and a peak of 58.15 with the full dataset. This tendency suggests the potential for further refinement of the filtering method. Moreover, to reduce computational cost, we use 5K items as the default scale in subsequent experiments.
Which generator and verifier LMMs should we use?
We next vary the capacity and identity of the generator and verifier LMMs used during data synthesis (Table 4). For the 3B student, pairing a GLM‑4.5V‑108B generator with a Qwen2.5‑VL‑72B verifier yields the best average 55.85; other high‑capacity pairs are close. For the 7B student, the same open‑weight verifier gives 57.56 with a GLM‑108B generator, while using GLM‑108B as both generator and verifier further nudges the average to 58.08. We keep the Qwen2.5‑VL‑72B verifier for the main results to maximize reproducibility with open weights, but Table 4 indicates that stronger verifier capacity translates to higher downstream accuracy.
Training approach and data source ablation.
Table 5 compares SFT vs RL from verification reward (RL) across three data sources: PMC-VQA (image–text pairs) (Zhang et al., 2023b), m23k (text‑only) (Huang et al., 2025a), and MedSynVQA. 1) RL outperform SFT for both 3B and 7B models, across all data source. 2) Under RL the MedSynVQA signal is the strongest, giving the best average on both 3B (55.85) and 7B (57.56). The results indicate that rubric‑based context-aware MedSynVQA dataset are more effective training source than the previous synthetic PMC-VQA (Zhang et al., 2023b) and the text-only one (Huang et al., 2025b).
Comparisons.
Table 6 summarizes head‑to‑head results on the full benchmark suite. Our students trained with MedSynVQA achieve 55.85 (3B) and 58.15 (7B), state‑of‑the‑art averages among open‑weight models considered. Notably, our 3B student surpasses MedVLThinker‑7B by +0.97 and all other 3–7B baselines; the 7B student improves over the best prior MedVLThinker‑7B by +3.27. Gains are consistent across datasets, with strong results on VQA‑RAD (up to 77.57).
Case study.
Figure 4 presents two cases, revealing deep comprehension with context for our generator and the leakage rejection by our verifier. Please refer to the appendix for more details.
Contamination analysis.
We practice contamination analysis between MedSynVQA and the evaluation suites in the appendix. No overlaps were found under this protocol.
5 Conclusions
MedVLSynther shows that high‑quality, auditable medical VQA data can be synthesized at scale from open biomedical literature by pairing rubric‑guided, context‑aware generation with a multi‑stage verifier. The resulting MedSynVQA delivers consistent gains for open‑weight LMMs across six benchmarks and ablations confirm that both the generator and verifier are necessary. Operating entirely on open data and models, the approach offers a reproducible, privacy‑preserving, and transparent path to supervision for medical VQA.

### v2

@@
 (2023b)’s 227k auto‑generated pairs (largely from text‑only LLMs) can omit modality cues, produce ambiguous stems, and yield visually ungrounded answers; broader compilations like Li et al.
-(2024) are closed, while modality‑specific (Hu et al., ) sets remain narrow.
+(2024) are closed, while modality‑specific (Hu and others, ) sets remain narrow.
 These limitations motivate a quality‑first strategy: we condition on figures, captions, and in‑text references and enforce a rubric‑guided generator plus a multi‑stage verifier to filter low‑quality items, yielding reliable, open data suitable for training medical LMMs without relying on private images.
@@
 Optional: localization or quantitative details when clearly supported.
-A small set of question archetypes (i.e., finding identification, diagnosis, next step, localization, modality recognition) reduces prompt entropy and encourages clinically meaningful questions.
+A small set of question archetypes (i.e., Finding/Abnormality Identification, Modality Recognition, Anatomy/Localization, Other Biological/Technical Attributes, Disease Diagnosis, Next Step, and Lesion Grading) reduces prompt entropy and encourages clinically meaningful questions.
 3.3 Multi‑stage, rubric‑based, context‑aware verification
@@
 | Qwen2.5-VL-3B-Instruct | 44.12 | 20.69 | 61.96 | 44.77 | 61.30 | 62.01 | 49.14 |
-| PMC-Style Text-only Generation | 48.82 | 20.40 | 63.38 | 51.08 | 73.08 | 72.06 | 54.80 |
+| PMC-Style Text-only Generation | 46.47 | 20.8 | 62.43 | 51.03 | 73.08 | 71.69 | 54.25 |
+| PMC-Style Text-Image Generation | 48.82 | 20.40 | 63.38 | 51.08 | 73.08 | 72.06 | 54.80 |
 | Rubric Context-Aware Generation | 52.35 | 20.60 | 62.49 | 51.83 | 70.43 | 70.59 | 54.72 |
@@
 | Qwen2.5-VL-7B-Instruct | 52.94 | 18.89 | 65.39 | 49.30 | 65.71 | 68.75 | 53.50 |
-| PMC-Style Text-only Generation | 51.76 | 21.70 | 64.31 | 53.43 | 68.03 | 71.69 | 55.15 |
+| PMC-Style Text-only Generation | 54.71 | 22.15 | 63.89 | 53.23 | 67.07 | 65.44 | 54.41 |
+| PMC-Style Text-Image Generation | 51.76 | 21.70 | 64.31 | 53.43 | 68.03 | 71.69 | 55.15 |
 | Rubric Context-Aware Generation | 58.24 | 23.50 | 65.41 | 53.83 | 68.03 | 75.00 | 57.33 |
@@
 Operating entirely on open data and models, the approach offers a reproducible, privacy‑preserving, and transparent path to supervision for medical VQA.
+Acknowledgments
+This work was partially funded by an unrestricted gift from Google.
+Appendix
+Appendix A The Use of Large Language Models (LLMs)
+During the preparation of this manuscript, we used OpenAI’s GPT-5 model for minor language refinement and smoothing of the writing.
+The AI tool was not used for generating original content, conducting data analysis, or formulating core scientific ideas.
+All conceptual development, experimentation, and interpretation were conducted independently without reliance on AI tools.
+Appendix B Reproducibility Statement
+We provide detailed instruction to reproduce our research.
+The prompts for generation and verification are presented in Section E.
+Section 3 provides instructions for initial data curation and filtering, question generation and verification, and our implementation details.
+Section 4.1 demonstrates the experimental settings, including benchmarks and comparison baselines.
+All models are trained on 8 A100 GPU machines using mixed precision.
+The hyperparameters are following “MedVLThinker” (Huang et al., 2025b).
+We full-stack release our data, code, and models in the pursuit of open science 111https://ucsc-vlaa.github.io/MedVLSynther/.
+Appendix C Ethical Considerations and Licensing
+MedVLSynther operates solely on PMC‑OA literature.
+We include source metadata and original licenses in dataset manifests and exclude items whose licenses are incompatible with redistribution for research.
+We provide attribution to source articles in per‑item metadata.
+MedVLSynther is intended exclusively for research; no clinical use is authorized.
+We release prompts, rubric, and filtering logs to enable community auditing.
+Appendix D More Data Statistics
+| Image Secondary Label | # Images | Ratio | Image Secondary Label | # Images | Ratio |
+| clinical imaging | 1039 | 7.17 | intraoral imaging | 393 | 2.71 |
+| brain | 984 | 6.79 | eye | 379 | 2.62 |
+| x-ray radiography | 879 | 6.07 | electrocardiography | 377 | 2.60 |
+| light microscopy | 829 | 5.72 | mammography | 374 | 2.58 |
+| computerized tomography | 785 | 5.42 | patient photo | 357 | 2.46 |
+| immunohistochemistry | 676 | 4.67 | procedural image | 335 | 2.31 |
+| magnetic resonance | 673 | 4.65 | laryngoscopy | 314 | 2.17 |
+| surgical procedure | 565 | 3.90 | microscopy | 65 | 0.45 |
+| skin lesion | 563 | 3.89 | scientific illustration | 31 | 0.21 |
+| functional magnetic resonance | 525 | 3.62 | line plot | 27 | 0.19 |
+| angiography | 519 | 3.58 | bar plot | 23 | 0.16 |
+| intraoperative image | 504 | 3.48 | immunoblot | 12 | 0.08 |
+| specimen | 500 | 3.45 | confocal microscopy | 11 | 0.08 |
+| optical coherence tomography | 492 | 3.40 | aerial photography | 2 | 0.01 |
+| ultrasound | 486 | 3.35 | table | 2 | 0.01 |
+| endoscopy | 477 | 3.29 | tree | 1 | 0.01 |
+| teeth | 470 | 3.24 | signal plot | 1 | 0.01 |
+| skull | 415 | 2.86 | user interface | 1 | 0.01 |
+| plot | 400 | 2.76 | ambiguous | 1 | 0.01 |
+Appendix E Prompts
+Appendix F Data Contamination Analysis
+Text decontamination.
+We audit overlap between the training pool (13,087 rows) and the held‑out test set (8,220 rows) over the fields question and options, after lowercasing, whitespace cleanup, digit masking (all numbers to “<NUM>”), and normalizing options to “A./B./…”.
+1) MinHash + Levenshtein hashes 3‑character n‑grams and confirms candidates with normalized edit similarity (, following Pal et al.
+(2022)).
+This conservative pass finds no near‑duplicate strings (pairs=0; hit queries=0; hit rate=0.0).
+2) Embedding + FAISS (Douze et al., 2024) encodes text with BAAI/bge‑m3 (Chen et al., 2024a) (L2‑normalized), retrieves top‑k=5 train neighbors per test item, and flags pairs with cosine larger then 0.88 (Tumre et al., 2025).
+This yields 47 pairs across 23 test queries (hit_rate about 0.280%), with per‑query hits mean=2.0 (max=5).
+The top‑1 similarity for those 23 queries averages 0.899 (median 0.894; p95=0.926), and qualifying pair similarities concentrate in [0.88, 0.95).
+Considering all test items, the best‑neighbor (“MaxSim”) distribution is mean 0.651 (median 0.644; p95=0.778), and Overlap@0.88 = 23/8,220 = 0.280%, indicating minimal distributional leakage beyond trivial lexical matches.
+We found no overlap of texts.
+Image decontamination.
+We repeat the two‑pass audit on images decoded directly from the dataset (using the first image per record by default).
+1) A hashing pipeline detects exact and near duplicates via MD5 of raw pixels and a 64‑bit perceptual hash (pHash); pHash neighbors are retrieved with a binary FAISS index, and we summarize each test image’s minimum Hamming distance and “Overlap@d” (fraction with best neighbor less than or equal to “d”).
+2) A semantic pipeline encodes images with OpenCLIP or BiomedCLIP (Ilharco et al., 2021; Zhang et al., 2023a), searches the train set with FAISS (Douze et al., 2024) (inner product on L2‑normalized features), and reports per‑test MaxSim, Overlap@, and a histogram of high‑similarity test–train pairs (e.g., at ).
+Together, hashing probes pixel-level or layout‑level duplication while embeddings probe semantic or style reuse; all statistics are computed from the images as stored in parquet under this single‑image‑per‑record protocol.
+We found no overlap of images.
+Appendix G Expert Evaluation Statistics
+To complement the downstream results and decontamination analysis, we conducted a small-scale human evaluation of MedSynVQA questions to assess intrinsic item quality and the behavior of the verifier.
+Five board-certified imaging experts participated, with clinical backgrounds covering oncologic CT/MR, abdominal and gastrointestinal CT/MR, brain CT/MR, diagnostic radiology, and gastrointestinal endoscopy.
+Some questions were jointly reviewed by more than one expert when they spanned multiple organs or modalities.
+G.1 Verifier-accepted items
+We first assessed the precision of the verifier on items it accepts.
+We randomly sampled 15 multiple-choice questions (MCQs) that had been accepted under the main verification threshold.
+Each item was rated on four 1 to 4 Likert scales (4 = excellent, 1 = poor): medical correctness and uniqueness of the keyed answer, clarity and professional wording, image grounding, and option design.
+In addition, experts provided a binary judgment on whether the item was acceptable for use as a medical MCQ.
+Table 8 summarizes the aggregate statistics, and 10 reports per-item ratings.
+14 of the 15 items (about 93%) were judged acceptable overall; one item was rejected because the brain MR image had such poor contrast that the expert considered the image non-evaluable, independent of the question text.
+The mean scores across criteria were 3.35 for medical correctness, 3.13 for clarity and professional wording, 3.46 for image grounding, and 3.40 for option design, corresponding to good to excellent quality on average.
+Two items received a score of 1 on the clarity and wording scale, indicating that while their medical content and image use were sound, the phrasing could be substantially improved.
+This is consistent with our finding that the current vocabulary related checks in the verifier do not fully capture wording quality and suggests that prompt refinement for clarity is a promising direction.
+G.2 Verifier-rejected items
+We then examined whether the verifier is justified when rejecting items.
+We randomly sampled 20 MCQs that had been filtered out by the verifier and, for each item, selected one essential rubric that had received a zero score.
+The audited rubrics and their frequencies, along with human agreement, are summarized in Table 9; Table 11 provides per-item details.
+Across the 20 rejected items, the audited essentials were: stem self-contained (5 items), vocabulary constraint (4), diagnosis leak (1), single correct option (2), option type consistency (4), clinical validity (2), and image–text consistency (2).
+For rubric types that do not require strong medical expertise (stem self-contained, vocabulary constraint, option type consistency), the authors first inspected the items.
+Stems marked as non self-contained were indeed unusable without external context, for example explicitly referring to “according to the context” rather than being answerable from the stem and image alone.
+Option type consistency failures involved heterogeneous option formats or options literally written as “none” (for example, option E is just “none” as an empty placeholder, rather than a meaningful choice such as “no abnormality” or “none of the above”), so these rejections were appropriate.
+For the vocabulary constraint rubric, we identified a systematic issue.
+In the detailed cases examined, the verifier incorrectly flagged sex, age, or clinical history as “unsupported,” even though these attributes were present in the caption or context.
+These are genuine false negatives: the current rubric prompt does not reliably capture linguistic professionalism or terminology use and can misfire when the model fails to correctly read the context.
+We therefore regard vocabulary constraint as a weaker signal and a primary target for future prompt and rubric refinement.
+For the medically heavy essentials (diagnosis leak, single correct option, clinical validity, and image-text consistency), items were evaluated by clinicians.
+In all inspected cases, experts agreed with the verifier.
+Items rejected for diagnosis leak did reveal the diagnosis directly in the stem; items rejected for single correct option or clinical validity had multiple plausible keys or medically incorrect statements; items rejected for image-text consistency did not match the visual evidence.
+G.3 Summary
+Although this human study is necessarily small due to limited clinician time and the difficulty of reviewing full image-based MCQs, it already involves five specialists and includes items that required cross-domain input.
+The audit shows that a large majority of verifier-accepted items are judged medically sound and usable, and that the essential medical rubrics driving rejection (diagnosis leak, single correct option, clinical validity, image–text consistency) align well with clinician judgment.
+The main systematic discrepancy is concentrated in the vocabulary-related check, which we now explicitly highlight as a limitation.
+Overall, Tables 8–11 provide intrinsic evidence about the reliability and failure modes of the verification stage and complement the downstream metrics and decontamination analyses in the main paper.
+| Subset | Items | Experts | Human pass rate | Mean medical correctness | Mean clarity | Mean image grounding | Mean option design |
+| Verifier-accepted MCQs | 15.00 | 5.00 | 14 / 15 (93%) | 3.35 | 3.13 | 3.46 | 3.40 |
+| Verifier-rejected MCQs | 20.00 | 5.00 | N/A | N/A | N/A | N/A | N/A |
+| Essential rubric | Count | Human agreement with verifier | Notes |
+| Stem self-contained | 5 | Yes (all) | Stems depended on external context, not answerable from stem+image |
+| Vocabulary constraint | 4 | No (in 3 cases) | Verifier wrongly flagged sex/age/history as unsupported |
+| Diagnosis leak | 1 | Yes (all) | Diagnosis explicitly revealed in stem |
+| Single correct option | 2 | Yes (all) | Multiple plausible keys or ambiguous answer |
+| Option type consistency | 4 | Yes (all) | Heterogeneous option formats or meaningless “none” placeholder |
+| Clinical validity | 2 | Yes (all) | Medically incorrect or inconsistent statements |
+| Image–text consistency | 2 | Yes (all) | Text description did not match the visual findings |
+| Item ID | Reviewer specialty | Human acceptable? | Medical correctness (1–4) | Clarity & wording (1–4) | Image grounding (1–4) | Option design (1–4) | Notes (optional) |
+| A01 | diagnostic radiology | Yes | 3 | 3 | 4 | 3 | |
+| A02 | diagnostic radiology | Yes | 4 | 3 | 4 | 3 | |
+| A03 | diagnostic radiology | Yes | 4 | 4 | 4 | 4 | |
+| A04 | diagnostic radiology | Yes | 4 | 4 | 4 | 4 | |
+| A05 | diagnostic radiology | Yes | 4 | 4 | 4 | 4 | |
+| A06 | diagnostic radiology | Yes | 3 | 3 | 3 | 3 | |
+| A07 | gastrointestinal endoscopy | Yes | 4 | 4 | 4 | 4 | |
+| A08 | gastrointestinal endoscopy | Yes | 3 | 3 | 3 | 3 | |
+| A09 | gastrointestinal endoscopy | Yes | 4 | 3 | 3 | 3 | |
+| A10 | gastrointestinal endoscopy | Yes | 2 | 3 | 2 | 3 | |
+| A11 | brain CT/MR | Yes | 4 | 4 | 4 | 4 | |
+| A12 | brain CT/MR | No | N/A | 1 | 3 | 3 | poor contrast images |
+| A13 | brain CT/MR | Yes | 4 | 4 | 4 | 4 | |
+| A14 | brain CT/MR | Yes | 2 | 3 | 2 | 4 | |
+| A15 | brain CT/MR | Yes | 2 | 1 | 4 | 2 |
+| Item ID | Reviewer specialty | Essential rubric checked | Human agrees with rejection? |
+| B01 | author self-review | Stem self-contained | Yes |
+| B02 | author self-review | Stem self-contained | Yes |
+| B03 | author self-review | Stem self-contained | Yes |
+| B04 | author self-review | Stem self-contained | Yes |
+| B05 | author self-review | Stem self-contained | Yes |
+| B06 | author self-review | vocabulary constraint | Yes |
+| B07 | author self-review | vocabulary constraint | No |
+| B08 | author self-review | vocabulary constraint | No |
+| B09 | author self-review | vocabulary constraint | No |
+| B10 | oncologic CT/MR, diagnostic radiology | diagnosis leak | Yes |
+| B11 | oncologic CT/MR | single correct option | Yes |
+| B12 | gastrointestinal CT/MR, gastrointestinal endoscopy | single correct option | Yes |
+| B13 | author self-review | option type consistency | Yes |
+| B14 | author self-review | option type consistency | Yes |
+| B15 | author self-review | option type consistency | Yes |
+| B16 | author self-review | option type consistency | Yes |
+| B17 | gastrointestinal CT/MR, gastrointestinal endoscopy | clinical validity | Yes |
+| B18 | gastrointestinal CT/MR, gastrointestinal endoscopy | clinical validity | Yes |
+| B19 | gastrointestinal CT/MR, gastrointestinal endoscopy | image-text consistency | Yes |
+| B20 | brain CT/MR | image-text consistency | Yes |